OpenStudy (anonymous):

Solve the equation. (1 over 3x + 9) – (2 over x + 3) = 2 A.x = 1 B.x = –23/6 C.x = –5/2 D.x = –5 I got -1/9 and that's not one of the answers. What did I do wrong? What's the corrects answer?

OpenStudy (anonymous):

\[\frac{ 1 }{ 3x+9 } - \frac{ 2 }{x+3 } =2\]

OpenStudy (anonymous):

\[\frac{ 1 }{3 (x+3) } - \frac{ 2 }{ x+3 } = 2\]

OpenStudy (anonymous):